Circular No. 6334

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) R. M. West, European Southern Observatory (ESO), reports that an

image of the complex tail system of C/1996 B2 was obtained by O. Pizarro with the ESO 1-m Schmidt telescope at La Silla on Feb. 28.36 UT under good observing conditions. The following is a description from an amplified reproduction of the 60-min Kodak Pan 4415 film (+ GG385 filter; 390–700 nm) by H.-H. Heyer (ESO, Garching): “A broad and featureless dust tail is seen at p.a. 280^{o} until 40′ from the nucleus. A narrow ion tail extends to 20′ at p.a. 289^{o}, where it branches into two main components — a broader one at p.a. 289^{o} with complex structure (and strong kinks at 35′ and 96′) and total length exceeding 160′ (plate edge), and a narrow one with sharp borders between p.a. 300^{o} and 301^{o} to ∼ 100′ distance. There are several streamers in the interval p.a. 305^{o}–315^{o}, some reaching to ∼ 25′.”

Total visual magnitude and coma-diameter estimates: Mar. 1.77 UT, 6.2, – (R. H. McNaught, Siding Spring Observatory, naked eye); 2.53, 5.8, 15′ (C. S. Morris, Whitaker Peak, CA, 10×50 binoculars; 1^{o}.75 tail in p.a. 278^{o}); 3.78, 5.7, – (G. Garradd, Loomberah, Australia, naked eye); 4.24, 5.5, 13′ (A. Pereira, Cabo da Roca, Portugal, 9×34 binoculars).

GRO J1744–28 M. J. Stark, University of Maryland at College Park, and the Labora-

tory for High Energy Astrophysics, Goddard Space Flight Center (LHEA/ GSFC); A. Baykal, National Research Council and LHEA/GSFC; T. Strohmayer, Universities Space Research Association and LHEA/GSFC; and J. H. Swank, LHEA/GSFC, write: “We report a correction to IAUC 6324 and confirmation with XTE of the result reported by Jung et al. on IAUC 6321. On IAUC 6324, we reported an incorrect direction of the phase shift in the pulsar arrival times of GRO J1744–28. Following each of the x-ray bursts, we reported as an advance in the pulse arrival time what is actually a lag. We reported the correct magnitude (25 ms) and duration (onset < 25 s; decay ∼ 1500 s). Combining the data from seven bursts observed on Jan. 18, we have confirmed a lag during the bursts in pulse-arrival times of 15 percent. This lag persists for the duration of the burst (8–10 s) in the majority of pulses, which keep the same shape. For a few periods at high intensity, the pulse is not present or is doubly peaked. After the burst, the pulse-arrival time lag drops to 5 percent within a few seconds and decays further in ∼ 1500 s, as we reported earlier.”

1996 March 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green